Simplify both sides as much as possible, and then solve:
18-13=1/3a+1/2a-5/12a

Answer by ewatrrr(24785)


Hi
Solving the following for a:
18-13 =1/3a + 1/2a -5/12a
5 =4/12a + 6/12a -5/12a
5 = 5/12a
60/5 = a
12 = a